OpenStudy (anonymous):

Combustion analysis of an unknown compound provides the following data: 73.5 grams C, 4.20 grams H, and 72.3 grams Cl. What is the percent composition of each element in this compound? A) 73.5% C, 4.20% H, 72.3% Cl B) 48.0% C, 3.00% H, 50.0% Cl C) 17.5% C, 1.00% H, 17.2% Cl D) 49.0% C, 2.80% H, 48.2% Cl

OpenStudy (matt6288):

all it is asking is for mass %. Add the masses of all the elements, and find the percentages by mass

OpenStudy (matt6288):

You have 73.5 grams C, 4.20 grams H, and 72.3 grams Cl added together it is 150 g so composition is: C 73.5/150 = 0.49 x 100% = 49% H 4,20/150 = 0.028 x 100% = 2.8 % Cl 72,3/150 = 0.482 x 100% = 48.2 % It's D